Write the equation of the line passing through (2,0) and (0,-5)
Answer by vertciel(183)

First, we need to find the slope.
Slope = change in y/change in x
= -5 - 0/0 - 2
= -5/-2
= 2.5 or just 5/2

Equation of the line: y - y1 = m(x - x1)
y - y1 = 5/2(x - x1)
Take one of the two points and substitute it into the equation. In this case, I'll pick (2,0) because it is easier to work with; no negatives.
y - 0 = 5/2(x - 2)
y = 5/2x - 5
or 5/2x - y - 5 = 0